OpenStudy (anonymous):

lim as x->-infin. of (X^5)(e^x) Use L'hospital's Rule.

OpenStudy (mathteacher1729):

First of all, which indeterminate form do we get here? 0/0, infinity / infinity, 0^0, infinity * infinity, etc.?

OpenStudy (anonymous):

infinity * infinity

OpenStudy (anonymous):

Not quite. You have -inf*(1/inf)

OpenStudy (anonymous):

oh ok... then what? that's not one of the forms to do L'hospital's rule. so do I have to do implicit differentiation?

OpenStudy (anonymous):

It is technically infinity*0 I believe is indeterminate. Rewrite it as a quotient: \[\lim_{x \rightarrow -\infty}\frac{x^5}{e^{-x}}\] Differentiate top and bottom until you get rid of the x and you get: \[\lim_{x \rightarrow -\infty}\frac{\frac{d^5}{dx^5}x^5}{\frac{d^5}{dx^5}e^{-x}}=\lim_{x \rightarrow -\infty}\frac{120}{-e^{-x}}=0\]

OpenStudy (anonymous):

If I see it right.

myininaya (myininaya):

\[\frac{x^5}{\frac{1}{e^x}}=\frac{x^5}{e^{-x}}\] so e^{-x}->inf as x->-inf x^5->-inf as x->-inf so we have \[\lim_{x \rightarrow -\infty}\frac{5x^4}{-e^{-x}}=\lim_{x \rightarrow -\infty}\frac{5x^4}{-e^{-x}}=\lim_{x \rightarrow -\infty}\frac{20x^3}{e^{-x}}=\lim_{x \rightarrow -\infty}\frac{60x^2}{-e^{-x}}\] =\[\lim_{x \rightarrow -\infty}\frac{120x}{-e^{-x}}=\lim_{x \rightarrow -\infty}\frac{120}{e^{-x}}=\lim_{x \rightarrow -\infty}120e^{-x}=120\lim_{x \rightarrow -\infty}e^{-x}=\infty\] let me check with graph

OpenStudy (anonymous):

myininaya you brought the e^-x up as a -x!!!! Its supposed to be positive :PPPP

myininaya (myininaya):

oh i see my last two steps oops

myininaya (myininaya):

\[\lim_{x \rightarrow -\infty}120e^x=120 \lim_{x \rightarrow -\infty}e^x=120(0)=0\]

myininaya (myininaya):

gj malevolence

OpenStudy (anonymous):

You too :P I skipped my differentiation xP

myininaya (myininaya):

i cant believe i did that

myininaya (myininaya):

i dont make mistakes in the words of satellite it was a type-0

myininaya (myininaya):

:)

OpenStudy (anonymous):

ok thank you.... why did you keep differentiating?

myininaya (myininaya):

because we had indeterminate form repeatedly

OpenStudy (anonymous):

You have to differentiate it until you get to a point where it is no longer indeterminate.

myininaya (myininaya):

we kept having either \[-\frac{\infty}{\infty} or \frac{\infty}{\infty}\] until we got to 120e^x

OpenStudy (anonymous):

ok I see it now... thank you :)